OpenStudy (anonymous):

Find the vertices, foci, and eccentricity of the ellipse. 56x^2 + 7y^2 = 8

OpenStudy (anonymous):

Hello, The first approach to Ellipse is to understand it's structure. \[\frac{ x^2 }{ a^2 }+\frac{ y^2 }{ b^2 }=1\] \[\frac{ y^2 }{ a^2 }+\frac{ x^2 }{ b^2 }=1\] center: (h,k) Focus/Foci: (±c,0) Vertex/Vertices: (±a,0)